Digital Test
Hewlett-Packard 524B

Manufacturer: Hewlett-Packard
Model: 524B
Year: 1957
Schematic: RE
Max. Rate: 10 MHz
Display: 1-of-10 neon, analog meters
Digits: 8
Logic: tube

The HP524 was a remarkable instrument in it's day with it's high count rate. This is the 'B' version. It utilises 84 vacuum tubes, physically is approximately a cube of 19-inch rack-width (~ 4 cubic feet), weighs in at 110 to 120 pounds and consumes 600 watts of power. In the 1950s this is what was required to count to just 10MHz!

The photo shows the unit displaying a count of "647852.84".

While this particular unit required days of cleaning after I received it, it had remarkably little wrong with it electronically. One tube, two capacitors and one resistor solved the electronic problems.

Note: To adjust units- and tens-digit meters: with MODE=MANUAL GATE, PERIOD UNIT=100KC, GATE=OPEN, adjust CENTER pots for both meters to bring needle to dot between 4 & 5. Set GATE=CLOSED, press RESET, now adjust ZERO pots.

Note: Fan must be operating when chassis is in cabinet. If not, thermal cutout may trip after a few minutes.

For more about the 524 models see the HP 520 Series.

The 524B with a later-model (different panel colour) 526B Time-Interval plug-in plugged in. The 524B was counting during the exposure of this photograph. Note the digit which has progressed from 7 thru 0 to 1, with the consequent carry incrementing the next digit from 2 to 3 (click photo for larger view).

Top view showing AC-4 counter string. The meter for the Tens digit is visible in the upper-left corner, along with pots for centering and zeroing the Units- and Tens-digit meters.

Rear deck: Power Supply on the right and Timebase on the left.

Right-side deck: Units & Tens Counters.

This is essentially the circuitry of the HP 520A Nuclear Scaler, the model which started the 520 Series. This also gives an idea of the complexity required to get the first decade to count at 10Mhz.

Bottom deck: Gate and Control.

Note the two massive power supply transformers at the bottom.

- Unit Log -

Serial Number: 2171
Year of Manufacture: 1957 (Capacitor stamped with "1-57", coils with "APR 57")
Other Identification: Case stamped with "BCE Co.", which would likely be the B.C. Electric Company.
Date of Receipt: 24 Mar 2008
State upon Receipt: Non-functional. Scratches in paint on top of case. Glass on crystal oven temperature indicator broken. Lower right snap pin for plug-in unit broken. Line cord chopped off. GATE lamp cover missing. Fairly grubby, will need good cleaning.

Date: 2008 Apr 02
Procedure: Cleaning in progress: tubes wiped off, tube clips washed; chassis, wiring and components wiped with damp sponge. Three control linkages dismantled to clean out dried grease.

Date: 2008 Apr 02
Procedure: V235 replaced with GE 5726 so hold-down clip will fit. Previous BRIMAR replacement was oversized bulb.

Date: 2008 Apr 05
Procedure: Minor soldering fixup of some wiring in the timebase section. C328 appears to have been previously replaced.

Date: 2008 Apr 07
Procedure: Broken plug-in slot snap-pin (lower-right) fixed by drilling and tapping pin for 4-40 screw. Broken side of 8-pin Amphenol socket for plug-in glued back on.

Date: 2008 Apr 07
Procedure: New line cord installed. Replacement GATE lamp cover installed. Darkened NE-51 STANDBY lamp replaced.

Date: 2008 Apr 08
Procedure: Powered up. Power supply voltages a little off but not much, adjustments tweaked. Counters will reset to 0, display cycle (as evidenced by gate lamp) functions, so timebase and gate circuitry are functioning, but unit does not count.

Date: 2008 Apr 08
Symptoms: On second power up, nothing functions properly, gate lamp does not go off completely.
Analysis: Voltage levels too high. 117V across 0B2, should be 108.
Solution: 0B2 replaced.

Date: 2008 Apr 09
Symptoms: In MANUAL GATE mode, unit will count on 100KC and 10MC but not on 10CPS or 1KC.
Analysis: 6.6V at pin 7 (grid input) of Trigger Unit, should be 0V. C209 (0.22 'bumble-bee' between V201.5 plate and Trigger Unit input) leaky.
Solution: C209 replaced.

Date: 2008 Apr 09
Procedure: Meter pots adjusted.

Date: 2008 Apr 09
Symptoms: 100KC Check results in 110.000 on 1S gate.
Solution: 1S phantastron adjusted.

Date: 2008 Apr 15
Symptoms: 100KC Check results in 110.0000 on 10S gate. 10S phantastron adjusted, but will not go low enough.
Solution: Original 2.7M resistor measures over 4M, replaced. Previously replaced 15M/0.5W resistor replaced with 18M/1W. Previously replaced C328 replaced.

Date: 2008 Apr 15
Procedure: Other alignments (GATE LENGTH, FREQ. TRIM, PERIOD TRIM) adjusted. Reassembled. Fully functional.

  .. | Hewlett-Packard 524B | ..
Early Digital Test Equipment